# Can EDA Tool Feedback Improve Verilog Generation by LLMs?

### v1

@@ Abstract.
Traditionally, digital hardware designs are written in the Verilog hardware description language (HDL) and debugged manually by engineers. This can be time-consuming and error-prone for complex designs. Large Language Models (LLMs) are emerging as a potential tool to help generate fully functioning HDL code, but most works have focused on generation in the single-shot capacity: i.e., run and evaluate, a process that does not leverage debugging and as such does not adequately reflect a realistic development process.
In this work we evaluate the ability of LLMs to leverage feedback from electronic design automation (EDA) tools to fix mistakes in their own generated Verilog. To accomplish this we present an open-source, highly customizable framework, AutoChip, which combines conversational LLMs with the output from Verilog compilers and simulations to iteratively generate and repair Verilog. To determine the success of these LLMs we leverage the VerilogEval benchmark set. We evaluate four state-of-the-art conversational LLMs, focusing on readily accessible commercial models.
EDA tool feedback proved to be consistently more effective than zero-shot prompting only with GPT-4o, the most computationally complex model we evaluated. In the best case we observed a 5.8% increase in the number of successful designs with a 34.2% decrease in cost over the best zero-shot results. Mixing smaller models with this larger model at the end of the feedback iterations resulted in equally as much success as with GPT-4o using feedback, but for an additional 41.9% less cost (overall decrease in cost over zero-shot of 89.6%).
@@ 1. Introduction
Designing digital hardware with a hardware description language (HDL), such as Verilog or VHDL, is a niche skill and part of a demanding process requiring substantial expertise. Any mishaps can lead to implementations fraught with bugs and errors (Dessouky et al., 2019), and with growing demand for digital systems, there is a growing demand for techniques that can assist in generating quality HDL. High-level synthesis (HLS) tools, for instance, are able to transform designs written in high-level software languages like C to target HDLs and implement digital hardware.
Recent efforts have shifted the abstraction level higher, leveraging state-of-the-art Large Language Models (LLMs) (Vaswani et al., 2017) to translate natural language to Verilog. DAVE (Pearce et al., 2020) and VeriGen (Thakur et al., 2023a) were the first efforts seeking to fine-tune LLMs specifically to generate Verilog. However, VeriGen and its ilk were investigated for their use in a zero-shot manner, i.e., they output code in response to a prompt. However, designing hardware in the real-world does not work this way—code is rarely correct on the first try. Instead, hardware designers iterate over their designs, using feedback from simulation and synthesis tools to identify and fix bugs so that an implementation will meet design specifications. This feedback-based approach is not well reflected in existing code-generation LLMs. Recent work (Blocklove et al., 2023) has proposed an iterative and interactive conversational (or chat-based) approach for Verilog code generation, more closely mimicking the design flow of a human hardware engineer. In this case, though, feedback comes entirely from a human developer who inspects the code, identifies bugs, and provides detailed feedback to the LLM. Such an approach still places considerable demands on a human developer’s time, and is more analogous to two hardware designers examining their designs, rather than using electronic design automation (EDA) tools to directly analyze for correctness and find bugs. We therefore ask: Can further automation reduce the burden on the designer?
@@
To help answer this question, we have developed and refined a framework for automating the hardware design process using LLMs, called AutoChip. While originally a simple iterative loop (Thakur et al., 2023b), as we discuss further in Section 3, we observed that the initial code generated from a prompt had a significant impact on the trajectory of the design flow and the eventual success or failure of the LLM-generated design. As such, we developed a more expansive tree search methodology (Figure 1), enabling us to more completely evaluate the ability of LLMs to use tool-based feedback to repair their own HDL in the same manner as a hardware engineer. Starting with a design prompt, AutoChip creates and evaluates $k$ candidate solutions and then enhances the most successful design by identifying and rectifying compilation errors and functional bugs over repeated interactions with an LLM. Each candidate design is analyzed for compilation errors/warnings and/or incorrect test cases from a testbench. These messages are used to rank the candidate solutions, and we return feedback from the tools and testbenches for the best candidate with a prompt to the LLM to refine its implementation and generate $k$ more candidates. This process is followed until all tests pass for a candidate response or a tree depth of $d$ iterations is reached, at which point the most successful candidate design from the tree search is returned.
AutoChip was evaluated with two feedback modes: ‘full context’ keeps appending prompts and responses to the ‘conversation’ with the LLM; and ‘succinct’ prompts only with feedback from the most recent iteration of the framework to try ensure that the process ‘fits’ within the context windows of LLMs.
@@
Our findings for these questions provide useful insights into how current LLMs can be leveraged to enable a fully automated chip design process, given a natural language design specification and resulting in a completed chip design ready for tapeout.
@@ 2. Background and Prior Work
LLMs are machine learning (ML) models built with transformers and are trained in a self-supervised manner on vast language data sets. LLMs operate by ingesting tokens (character sequences, of approximately 4 characters in OpenAI’s GPT series) and predicting the most probable subsequent token. The most powerful LLMs, e.g., ChatGPT (OpenAI, 2022), Bard (Pichai, 2023), and Code Llama (Google, 2023), boast hundreds of billions of parameters (Brown et al., 2020; Chen et al., 2021) and generalize to a broad range of tasks. Their accuracy is boosted via instruction tuning and reinforcement learning with human feedback (Ouyang et al., 2022), allowing the LLMs to more effectively understand and respond to user intentions. Prior work specialized LLMs for code generation. GitHub Copilot (GitHub, 2021) was an early LLM-based code completion engine.
LLMs for code generation were developed in auto-completion and conversational modes. DAVE (Pearce et al., 2020) was the first LLM (finetuned GPT-2) for Verilog generation. VeriGen (Thakur et al., 2023a) improved upon this work by expanding on the size of the model and size of the data sets. Chip-Chat (Blocklove et al., 2023) evaluated ChatGPT-4 to work with a hardware designer to generate a processor and the first fully AI-generated tapeout. RTLCoder (Liu et al., 2024) is another lightweight model for generating Verilog.
To evaluate model performance, a variety of benchmarks have been presented alongside further LLM developments, such as VerilogEval (Liu et al., 2023b) which evaluates LLMs’ abilities to write Verilog on benchmarks from HDLBits. Similarly, RTLLM (Lu et al., 2023) provides a further set of benchmarks. Other works have examined LLMs for hardware in tasks such as hardware bug repair (Ahmad et al., 2024) and generating SystemVerilog assertions (Kande et al., 2024).
@@ 3. AutoChip Design Framework
Figure 1 illustrates AutoChip’s functional design. The input to AutoChip is a natural language description of the desired functionality with a Verilog module definition (i.e. the I/O) and an accompanying testbench. In our evaluations we leverage the VerilogEval (Liu et al., 2023b) dataset for the source of these descriptions and testbenches, though AutoChip is not restricted to these benchmarks. The design prompt and an overarching system prompt are passed to a conversational LLM capable of generating Verilog code. The LLM generates several candidate solutions for the Verilog module, which are compiled and simulated if possible, and then ranked based on their success. Should the response not contain a Verilog module it is given a rank of $-2$, if it fails to compile it is given a rank of $-1$, if the compilation has warnings it is given a rank of $-0.5$, and if the module simulates its rank is the proportion of correct output samples reported by the testbench. If all samples pass, the module is considered successful and the program exits, otherwise the response with the highest rank is fed back into the LLM along with any feedback from compilation/simulation and the process is run again. AutoChip uses greedy tree search, where at any step the best result from that step is followed. This is in contrast to prior work which leverages feedback for design, such as Chip-Chat (Blocklove et al., 2023) which only uses a single candidate response for iteration.
@@
Prompting Strategy: Three prompt types are used in AutoChip: ‘system/context’ prompt, ‘design’ prompt, and ‘feedback’ prompt.  Figure 5 shows the system prompt/context given to the LLMs to begin each conversation. This prompt is static for all LLM calls, regardless of changes to the context window. Our response parser detects module and endmodule statements to extract Verilog modules when the system prompt is not rigidly followed.
@@
GPT-4o-Mini (OpenAI, 2024b) 128K \$0.15 \$0.60
GPT-4o (OpenAI, 2024a) 128K \$5.00 \$15.00
@@
Choice of Context Window: The quality of LLM responses depends on the conversation’s context window. As conversational LLMs have token limits, keeping all responses and feedback is often infeasible. The context window needs to shift during the automated run to keep only the information necessary for the next run, referred to as using ‘succinct’ feedback instead of ‘full-context,’ where all messages are used. With ‘succinct’ feedback, when an LLM is prompted to fix an issue, only the most recently generated module and its associated errors are given to the LLM. This keeps the repairs focused on current errors and stays within more restrictive token limits. Table 2 offers the context window shifting per-iteration.
@@
With ‘full-context’ feedback, the LLM input grows until a successful design is generated, maximum depth is reached in AutoChip, or the LLM token length is exceeded.
@@ 4.1. Benchmarking Prompts and Testbenches
Design Prompts: To evaluate AutoChip we leverage the dataset from VerilogEval (Liu et al., 2023b), which includes prompts and testbenches for a significant selection of problems from HDLBits (Wong, 2017), a site for Verilog practice with problems ranging in difficulty from simple Verilog syntax questions to more abstract sequential circuits and debugging. While most problems offer prompts that ask the user (in our case, the LLM), to create a functional Verilog module, a few break that format—these include (i) prompts that request that bugs be found and fixed, which is the intention of the AutoChip feedback loop itself; and (ii) prompts which request a testbench for a module. Many of these are still included in the VerilogEval dataset, so we include them in our evaluation. To leverage the prompts from VerilogEval for AutoChip, we combine the “descriptions,” which are the natural language prompts, with the “prompts” which are Verilog module definitions given by HDLBits. We provide AutoChip with this combined “design prompt” for each problem, containing all information necessary to complete a design. VerilogEval leaves out some problem categories from HDLBits. Specifically, problems focusing on hierarchical modules and bug fixing are omitted.
@@ 5.1. Single-Model Feedback Results
To determine if feedback from hardware verification tools improves the results, we first established a set of depth and candidate parameters for the feedback tree. We query the evaluated LLMs at depths of $d=\{0,1,5,10\}$ and with $k=\{1,5\}$ candidates. A depth of 0 corresponds to no feedback being given at all, equivalent to other works’ zero-shot analysis, for which we performed additional tests with $k=\{25,30,55\}$ candidates. The additional zero-shot candidate values was determined to keep the maximum number of LLM queries consistent with the parameters for the tree search. For example, both a test with $d=10,k=5$ and a test with $d=0,k=55$ have the same potential maximum number of LLM queries, though if a functioning design is found before the maximum potential, the test would still end early.
Prior to completing the more extensive tests of tool feedback-based design generation, we evaluate the generated designs for both ‘succinct’ and ‘full-context’ feedback (RQ4). We restrict this analysis to Claude 3 Haiku and GPT-3.5-Turbo, as those models are both relatively inexpensive and provide insight into the effect that the feedback context can have (given significantly different token limits). Following the exploration of context length with the two simpler models, we identified that providing the ‘full-context’ feedback resulted in similarly successful designs, as shown in Section 5.1, while requiring far fewer tokens over the course of longer tree searches. As a result, we completed the remainder of the tree search analysis with only ‘succinct’ feedback to reduce both complexity and cost.
@@
We can further examine these results from two perspectives: model effort, estimated by the average input and output tokens needed to generate a successful design, and model complexity, estimated by the \$ USD costs needed to generate a successful design.
@@
Figure 7. Generated circuit success rates for each evaluated model, shown as the percentage of generated circuits which pass all tests given the average number of tokens needed. Pareto points are plotted with dashed lines. Higher ‘accuracy’ at lower ‘average tokens’ is better. Data points represented by ‘x’ indicate results from zero-shot testing.
@@
Model Complexity: Figure 7(a) and Figure 7(a) show a similar analysis, but rather than giving success as a function of the tokens used, they report success as a function of the \$ USD cost at the time of publication. We see that, often, more complex models are both more successful and use fewer tokens to reach that level of success; however, these larger models are more computationally complex, so while they use fewer tokens across complete tests, they do far more with those tokens. This difference in complexity is reflected in, though we don’t know how for certain, the cost to use those models from OpenAI and Anthropic. While the relationship between model complexity and cost charged is unknown, it stands to reason that the more complex and capable models would cost more. The costs of the models evaluated in this paper are given in Section 3, and are used along with the token usage data to determine the average functional cost of generating a design for each benchmark. While GPT-4o had the highest rate of success and required the fewest average number of tokens for that success, the cost of these tokens far outstrips any of the smaller models tested. The cost of generation with tool feedback, though, was still significantly lower than with equivalent maximum potential candidates evaluated as zero-shot (RQ3).

### v3

# Automatically Improving LLM-based Verilog Generation using EDA Tool Feedback

@@ Abstract.
Traditionally, digital hardware designs are written in the Verilog hardware description language (HDL) and debugged manually by engineers. This can be time-consuming and error-prone for complex designs. Large Language Models (LLMs) are emerging as a potential tool to help generate fully functioning HDL code, but most works have focused on generation in the single-shot capacity: i.e., run and evaluate, a process that does not leverage debugging and, as such, does not adequately reflect a realistic development process. In this work, we evaluate the ability of LLMs to leverage feedback from electronic design automation (EDA) tools to fix mistakes in their own generated Verilog. To accomplish this, we present an open-source, highly customizable framework, AutoChip, which combines conversational LLMs with the output from Verilog compilers and simulations to iteratively generate and repair Verilog. To determine the success of these LLMs we leverage the VerilogEval benchmark set. We evaluate four state-of-the-art conversational LLMs, focusing on readily accessible commercial models. EDA tool feedback proved to be consistently more effective than zero-shot prompting only with GPT-4o, the most computationally complex model we evaluated. In the best case, we observed a 5.8% increase in the number of successful designs with a 34.2% decrease in cost over the best zero-shot results. Mixing smaller models with this larger model at the end of the feedback iterations resulted in equally as much success as with GPT-4o using feedback, but incurred 41.9% lower cost (corresponding to an overall decrease in cost over zero-shot by 89.6%).
@@ 1. Introduction
Designing digital hardware with a hardware description language (HDL), such as Verilog or VHDL, is a niche skill and part of a demanding process requiring substantial expertise. Any mishaps can lead to implementations fraught with bugs and errors (Dessouky et al., 2019), and with growing demand for digital systems, there is a growing demand for techniques that can assist in generating quality HDL code. High-level synthesis (HLS) tools, for instance, are able to transform designs written in high-level software languages like C to target HDLs and implement digital hardware.
Recent efforts have shifted the abstraction level higher, leveraging state-of-the-art Large Language Models (LLMs) (Vaswani et al., 2017) to translate natural language to Verilog. DAVE (Pearce et al., 2020) and VeriGen (Thakur et al., 2023a) were the first efforts seeking to fine-tune LLMs specifically to generate Verilog. However, VeriGen and its ilk were investigated for their use in a zero-shot manner, i.e., they examined output code in response to a single prompting. However, designing hardware in the real world does not work this way—code is rarely correct on the first try. Instead, hardware designers iterate over their designs, using feedback from simulation and synthesis tools to identify and fix bugs so that an implementation will meet design specifications. This feedback-based approach is not well reflected in existing code-generation LLMs. Recent work (Blocklove et al., 2023) has proposed an iterative and interactive, conversational (or chat-based) approach for Verilog code generation, more closely mimicking the design flow of a human hardware engineer. In this case, though, feedback comes entirely from a human developer who inspects the code, identifies bugs, and provides detailed feedback to the LLM. Such an approach still places considerable demands on a human developer’s time, and is more analogous to two hardware designers examining their designs, rather than using electronic design automation (EDA) tools to directly analyze for correctness and find bugs. We therefore ask: Can further automation reduce the burden on the designer?
@@
To help answer this question, we have developed and refined a framework for automating the hardware design process using LLMs, called AutoChip. While originally a simple iterative loop (Thakur et al., 2023b), as we discuss further in Section 3, we observed that the initial code generated from a prompt had a significant impact on the trajectory of the design flow and the eventual success or failure of the LLM-generated design. As such, we developed a more expansive tree search methodology (Figure 1), enabling us to more completely evaluate the ability of LLMs to use tool-based feedback to repair their own HDL code in the same manner as a hardware engineer. Starting with a design prompt, AutoChip creates and evaluates $k$ candidate solutions and then enhances the most successful design by identifying and rectifying compilation errors and functional bugs over repeated interactions with an LLM. Each candidate design is analyzed for compilation errors/warnings and/or incorrect test outputs from a testbench. These messages are used to rank the candidate solutions, and we return feedback from the tools and testbenches for the best candidate with a prompt to the LLM to refine its implementation and generate $k$ more candidates. This process continues until all tests pass for a candidate response or a tree depth of $d$ iterations is reached, at which point the most successful candidate from the tree search is returned.
AutoChip was evaluated with two feedback modes: “full context” keeps appending prompts and responses to the “conversation” with the LLM; and “succinct” prompts only with feedback from the most recent iteration of the framework to try ensure that the process “fits” within the context windows of LLMs.
@@
Our findings provide useful insights into how current LLMs can be leveraged to enable a fully automated chip design process, given a natural language specification and resulting in a completed chip design for tapeout.
@@ 2. Background and Prior Work
LLMs are machine learning (ML) models built with transformers and are trained in a self-supervised manner on vast language data sets. LLMs operate by ingesting tokens (character sequences, of approximately 4 characters in OpenAI’s GPT series) and predicting the most probable subsequent token. The most powerful LLMs, e.g., ChatGPT (OpenAI, 2022), Bard (Pichai, 2023), and Code Llama (noa, 2023), boast hundreds of billions of parameters (Brown et al., 2020; Chen et al., 2021) and generalize to a broad range of tasks. Their accuracy is boosted via instruction tuning and reinforcement learning with human feedback (Ouyang et al., 2022), allowing the LLMs to more effectively understand and respond to user intentions. Prior work specialized LLMs for code generation. GitHub Copilot (GitHub, 2021) was an early LLM-based code completion engine.
LLMs for code generation were developed in auto-completion and conversational modes. DAVE (Pearce et al., 2020) was the first LLM (fine-tuned GPT-2) for Verilog generation. VeriGen (Thakur et al., 2023a) improved upon this work by expanding on the size of the model and size of the data sets. Chip-Chat (Blocklove et al., 2023) evaluated ChatGPT-4 to work with a hardware designer to generate a processor and the first fully AI-generated tapeout. RTLCoder (Liu et al., 2024) is another lightweight model fine-tuned specifically for generating Verilog. There have also been some projects, such as BetterV (Pei et al., 2024) and CodeV (Zhao et al., 2024), which use multiple LLMs, some fine-tuned, and other pre-processed data to improve the generated Verilog through additional tasks such as summarization of components and discriminator-guided generation, respectively. VeriSeek (Wang et al., 2024) leverages reinforcement learning with golden code feedback as a method of finetuning another LLM for Verilog. LLMs have also been used in conjunction with a Monte Carlo tree-search (MCTS) to optimize generated designed for power, performance, and area efficiency, focusing on generating various sizes of adders, multipliers, and multiply-accumulate units (DeLorenzo et al., 2024).
To evaluate model performance, a variety of benchmarks have been presented alongside further LLM developments, such as VerilogEval (Liu et al., 2023b; Pinckney et al., 2024) which evaluates LLMs’ abilities to write Verilog on benchmarks from HDLBits. Similarly, RTLLM (Lu et al., 2023) provides a further set of benchmarks. Other works have examined LLMs for hardware in tasks such as hardware bug repair (Ahmad et al., 2024) and generating SystemVerilog assertions (Kande et al., 2024).
@@ 3. AutoChip Design Framework
The original design of AutoChip (Thakur et al., 2023b) used a single iterative path to improve upon generated Verilog. This structure appears in some other works as well, such as ChipGPT (Chang et al., 2023) and VeriAssist (Huang et al., 2024).
Figure 1 illustrates AutoChip’s expanded tree search functionality. The input to AutoChip is a natural language description of the desired circuit with a Verilog module definition (i.e. the I/O) and an accompanying testbench. In this work, AutoChip has the limitation that it does not allow for the generation of partial designs and it requires a functioning testbench for feedback purposes.
In our evaluations, we leverage the VerilogEval (Liu et al., 2023b) dataset for the source of these descriptions and testbenches, though AutoChip is not restricted to these benchmarks. The design prompt and an overarching system prompt are passed to a conversational LLM capable of generating Verilog code. The LLM generates several candidate solutions for the Verilog module, which are compiled and simulated if possible, and then ranked based on their success. Should the response not contain a Verilog module it is given a rank of $-2$, if it fails to compile it is given a rank of $-1$, if the compilation has warnings it is given a rank of $-0.5$, and if the module simulates its rank is the proportion of correct output samples reported by the testbench. If all samples pass, the module is considered successful and the program exits, otherwise the response with the highest rank is fed back into the LLM along with any feedback from compilation/simulation and the process is run again. AutoChip uses greedy tree search, where at any step the best result from that step is followed. This is in contrast to prior work which leverages feedback for design, such as Chip-Chat (Blocklove et al., 2023) which only uses a single candidate response for iteration.
@@
Prompting Strategy: Three prompt types are used in AutoChip: “system/context” prompt, “design” prompt, and “feedback” prompt.  Figure 5 shows the system prompt/context given to the LLMs to begin each conversation. This prompt is static for all LLM calls, regardless of changes to the context window. Our response parser detects module and endmodule statements to extract Verilog modules when the system prompt is not rigidly followed.
@@
GPT-4o-Mini (OpenAI, [n. d.]) 128K \$0.15 \$0.60
GPT-4o (OpenAI, 2024) 128K \$5.00 \$15.00
@@
Choice of Context Window: The quality of LLM responses depends on the conversation’s context window. As conversational LLMs have token limits, keeping all responses and feedback is often infeasible. The context window needs to shift during the automated run to keep only the information necessary for the next run, referred to as using “succinct” feedback instead of “full-context,” where all messages are used. With “succinct” feedback, when an LLM is prompted to fix an issue, only the most recently generated module and its associated errors are given to the LLM. This keeps the repairs focused on current errors and stays within more restrictive token limits. Table 2 offers the context window shifting per-iteration.
@@
With “full-context” feedback, the LLM input grows until a successful design is generated, maximum depth is reached in AutoChip, or the LLM token length is exceeded.
@@ 4.1. Benchmarking Prompts and Testbenches
Design Prompts: To evaluate AutoChip we leverage the dataset from VerilogEval (Liu et al., 2023b), which includes prompts and testbenches for a significant selection of problems from HDLBits (Wong, 2017), a site for Verilog practice with problems ranging in difficulty from simple Verilog syntax questions to more abstract sequential circuits and debugging. Recently, the VerilogEval benchmark set was updated with a new prompt setup and framework (Pinckney et al., 2024), now referred to as VerilogEval v2. As this update occurred after the experimentation discussed here was completed, this work uses the VerilogEval v1 benchmarks. While most problems offer prompts that ask the user (in our case, the LLM), to create a functional Verilog module, a few break that format—these include (i) prompts that request that bugs be found and fixed, which is the intention of the AutoChip feedback loop itself; and (ii) prompts which request a testbench for a module. Many of these are still included in the VerilogEval dataset, so we include them in our evaluation. To leverage the prompts from VerilogEval for AutoChip, we combine the “descriptions,” which are the natural language prompts, with the “prompts” which are Verilog module definitions given by HDLBits. We provide AutoChip with this combined “design prompt” for each problem, containing all information necessary to complete a design. VerilogEval leaves out some problem categories from HDLBits. Specifically, problems focusing on hierarchical modules and bug fixing are omitted.
@@ 5.1. Single-Model Feedback Results
To determine if feedback from hardware verification tools improves the results, we first established a set of depth and candidate parameters for the feedback tree. We query the evaluated LLMs at depths of $d=\{0,1,5,10\}$ and with $k=\{1,5\}$ candidates. A depth of 0 corresponds to no feedback being given at all, equivalent to other works’ zero-shot analysis, for which we performed additional tests with $k=\{25,30,55\}$ candidates. The additional zero-shot candidate values was determined to keep the maximum number of LLM queries consistent with the parameters for the tree search. For example, both a test with $d=10,k=5$ and a test with $d=0,k=55$ have the same potential maximum number of LLM queries, though if a functioning design is found before the maximum potential, the test would still end early. Figure 7 traces the greedy tree search for an example of a complete run of AutoChip. The code from this successful path is shown in Figure 8.
Figure 7. The successful tree search path for the Human benchmark of review2015_fsmshift with 5 response candidates (R_#) and a maximum depth of 10. The LLM used for this example was GPT-3.5-Turbo. Code for the most successful candidates at each depth is given in Figure 8. Full logs for each step, as well as all other outputs, can be found at https://zenodo.org/records/13864552.
1module top_module(
2 input clk,
3 input reset,
4 output shift_ena);
5reg [3:0] count;
6reg shift_ena_reg;
7wire shift_ena;
8always @(posedge clk or posedge reset) begin
9 if (reset) begin
10 count <= 4’b0;
11 shift_ena_reg <= 1’b1;
12 end else begin
13 count <= count + 1’b1;
14 if (count < 4’b1000) begin
15 shift_ena_reg <= 1’b1;
16 end else begin
17 shift_ena_reg <= 1’b0;
18 end
19 end
20end
21assign shift_ena = shift_ena_reg;
22endmodule
((a)) Module generated in iteration 0.
1module top_module(
2 input clk,
3 input reset,
4 output reg shift_ena);
5reg [3:0] count;
6always @(posedge clk or posedge reset) begin
7 if (reset) begin
8 count <= 4’b0;
9 shift_ena <= 1’b1;
10 end else begin
11 count <= count + 1’b1;
12 if (count < 4’b1000) begin
13 shift_ena <= 1’b1;
14 end else begin
15 shift_ena <= 1’b0;
16 end
17 end
18end
19endmodule
((b)) Module generated in iteration 1.
1module top_module(
2 input clk,
3 input reset,
4 output wire shift_ena);
5reg [3:0] count;
6reg enable;
7always @(posedge clk) begin
8 if (reset) begin
9 count <= 4’b0;
10 enable <= 1’b1;
11 end else begin
12 if (count < 4’b1000) begin
13 count <= count + 1’b1;
14 end else begin
15 count <= 4’b0;
16 enable <= 1’b0;
17 end
18 end
19end
20assign shift_ena = enable;
21endmodule
((c)) Module generated in iteration 2.
1module top_module(
2 input clk,
3 input reset,
4 output wire shift_ena);
5reg [1:0] count;
6reg enable;
7always @(posedge clk) begin
8 if (reset) begin
9 count <= 2’b0;
10 enable <= 1’b1;
11 end else begin
12 if (count < 2’b11) begin
13 count <= count + 1’b1;
14 end else begin
15 count <= 2’b0;
16 enable <= 1’b0;
17 end
18 end
19end
20assign shift_ena = enable;
21endmodule
((d)) Module generated in iteration 3.
Figure 8. The four Verilog modules generated across the AutoChip tree search highlighted in Figure 7.
Prior to completing the more extensive tests of tool feedback-based design generation, we evaluate the generated designs for both “succinct” and “full-context” feedback (RQ4). We restrict this analysis to Claude 3 Haiku and GPT-3.5-Turbo, as those models are both relatively inexpensive and provide insight into the effect that the feedback context can have (given significantly different token limits). Following the exploration of context length with the two simpler models, we identified that providing the “full-context” feedback resulted in similarly successful designs, as shown in Section 5.1, while requiring far fewer tokens over the course of longer tree searches. As a result, we completed the remainder of the tree search analysis with only “succinct” feedback to reduce both complexity and cost.
@@
We can further examine these results from two perspectives: model effort, estimated by the average input and output tokens needed to generate a successful design, and model complexity, estimated by the U.S. dollar (\$ USD) costs needed to generate a successful design at the time of experimentation.
@@
Figure 9. Generated circuit success rates for each evaluated model, shown as the percentage of generated circuits which pass all tests given the average number of tokens needed. Pareto points are plotted with dashed lines. Higher “accuracy” at lower “average tokens” is better. Data points represented by ‘x’ indicate results from zero-shot testing.
@@
Model Complexity: Figure 9(a) and Figure 9(a) show a similar analysis, but rather than giving success as a function of the tokens used, they report success as a function of the \$ USD cost at the time of publication. We see that, often, more complex models are both more successful and use fewer tokens to reach that level of success; however, these larger models are more computationally complex, so while they use fewer tokens across complete tests, they do far more with those tokens. This difference in complexity is likely reflected by the varying costs for using different models, as determined by each model vendor, though how exactly to quantify this remains an open challenge. While the relationship between model complexity and cost charged is unknown, it stands to reason that the more complex and capable models would cost more.
The costs of the models evaluated in this paper are given in Section 3, and are used along with the token usage data to determine the average functional cost of generating a design for each benchmark. While GPT-4o had the highest rate of success and required the fewest average number of tokens for that success, the cost of these tokens far outstrips any of the smaller models tested. The cost of generation with tool feedback, though, was still significantly lower than with equivalent maximum potential candidates evaluated as zero-shot (RQ3).